# StreamDiffusion: A Pipeline-level Solution for Real-time Interactive Generation

### v1

@@ 3.4 Stochastic Similarity Filter
Note: We emphasize that compared to determine whether we skip the compute via a hard threshold, the proposed probability-sampling-based similarity filtering strategy leads to a more smooth video generation. Because the hard threshold is prone to making the video stuck which hurts the impression of watching video streaming, while the sampling-based method significantly improves the smoothness.
3.5 Pre-computation
The U-Net architecture requires both input latent variables and conditioning embeddings. Typically, the conditioning embedding is derived from a prompt embedding, which remains constant across different frames. To optimize this, we pre-compute the prompt embedding and store it in a cache. In interactive or streaming mode, this pre-computed prompt embedding cache is recalled. Within U-Net, the Key and Value are computed based on this pre-computed prompt embedding for each frame. We have modified the U-Net to store these Key and Value pairs, allowing them to be reused. Whenever the input prompt is updated, we recompute and update these Key and Value pairs inside U-Net.
For consistent input frames across different timesteps and to improve computational efficiency, we pre-sample Gaussian noise for each denoising step and store it in the cache. This ensures that while each denoising step has distinct noise, every timestep retains the same noise ($x_{t,\tau}\neq x_{t,\tau+1}$ and $x_{t+1,\tau}=x_{t,\tau}$). This approach is particularly relevant for image-to-image tasks.
$x_{t}=\sqrt{\alpha_{\tau}}x_{0}+\sqrt{\beta_{\tau}}\epsilon$ (12)
This is a minor point in low throughput scenarios, but at frame rates higher than 60 FPS, the overhead of recomputing these static values becomes noticeable. If we achieve 60 FPS (i.e. 16.666 ms) with recomputing these static values in the loop with pre-computed cache, we can achieve 62 FPS (i.e. 16.066 ms).
We note that we have a specific design of key-value cache for Latent Consistency Models (LCM). As per the original paper, we need to compute $c_{\mathrm{skip}}(\tau)$ and $c_{\mathrm{out}}(\tau)$ to satisfy the following equation:
3.6 Model Acceleration and Tiny AutoEncoder
Figure 6: GPU Usage comparison under static scene. (GPU: RTX3060, Number of frames: 20) The blue line represents the GPU usage with SSF, the orange line indicates GPU usage without SSF, and the red line denotes the Skip probability calculated based on the cosine similarity between input frames. Additionally, the top of the plot displays input images corresponding to the same timestamps. In this case, the character in the input images is only blinking. Thus, this analysis compares GPU usage in a static scenario.
@@ 4 Experiments
Table 1: Comparison of Average Inference Time at Different Denoising Steps with Speedup Factors. (GPU: NVIDIA RTX4090, CPU: Core i9-13900K, OS: Ubuntu22.04.3 LTS)
Denoising Step StreamDiffusion (ms) StreamDiffusion w/o TRT (ms) AutoPipeline Image2Image (ms)
@@
Table 2: Comparison of Average Inference Time at Different Denoising Steps among Different CFG Methods
Denoising Step Self-Negative RCFG (ms) Onetime-Negative RCFG (ms) Normal CFG (ms)
@@
Figure 9: Average inference time comparison between Stream Batch and normal sequential denoising with using TensorRT
Table 3: The effect of different modules on average inference time at different denoising steps. The different module ablations are validated on image-to-image generation.
Method Denoising Step Latency [ms]
Ours (txt2img) 1 9.42
4 26.30
Ours (img2img) 1 10.65
4 26.93
w/o stream batch 1 10.70
4 36.54
w/o TRT 1 21.34
4 48.15
w/o Pre-computation 1 13.63
4 31.16
w/o TinyAutoEncoder 1 49.72
4 65.48
w/o IO queue 1 15.13
4 29.72
w/o any optimization 1 76.78
4 134.37
with SSF 1 10.66
4 27.02
We implement StreamDiffusion pipeline upon LCM, LCM-LoRA [17, 18] and SD-turbo [21]. As a model accelerator, we use TensorRT and for the light weight efficient VAE we use TAESD [11]. Our pipeline is compatible to the consumer GPU. We test our pipeline on NVIDIA RTX4090 GPU, Intel Core i9-13900K CPU, Ubuntu22.04.3 LTS, and NVIDIA RTX3060 GPU, Intel Core i7-12700K, Windows11 for image generation. We note that we evaluate the throughput mainly via the average inference time per image through processing 100 images.
4.1 Quantitative Evaluation of Denoising Batch
The efficiency comparison between the denoising batch and the original sequential U-Net loop is shown in Fig. 8. When implementing a denoising batch strategy, we observe a significant improvement in processing time. It achieves a reduction by half when compared to a conventional U-Net loop at sequential denoising steps. Even though applying TensorRT, the accelerate tool for neural modules, our proposed Stream Batch still boosts the efficiency of original sequential diffusion pipeline by a large margin at different denoising steps.
Figure 10: Results using no CFG, standard CFG, and RCFG with Self-Negative and Onetime-Negative approaches. When compared to cases where CFG is not utilized, the cases with CFG utilized can intensify the impact of prompts. In the proposed method RCFG, a more pronounced influence of prompts was observed. Both CFG and RCFG use guidance scale $\gamma=1.4$. For RCFG, the first two rows use magnitude modelation coefficient $\delta=1.0$, and the third row uses $\delta=0.5$.
Besides, we compare our method with the AutoPipelineForImage2Image, which is a pipeline developed by Huggingface diffusers  [^1] https://github.com/huggingface/diffusers. The average inference time comparison is presented in Table. 1. Our pipeline demonstrates a substantial speed increase. When we use TensorRT, StreamDiffusion achieves a minimum speed-up 13.0 times when running the 10 denoising steps, and reaching up to 59.6 times in scenarios when involving a single denoising step. Even though without TensorRT, StreamDiffusion achieves 29.7 times speed up compared to AutoPipeline when using one step denoising, and 8.3 times speedup at 10 step denoising.
Table. 2 presents a comparison of the inference times for StreamDiffusion pipelines with RCFG and conventional CFG. The additional computations required to apply Self-Negative RCFG are merely lightweight vector operations, resulting in negligible changes in inference time compared to when Self-Negative is not used. When employing Onetime-Negative RCFG, additional UNet computations are necessary for the first step of the denoising process. Therefore, One-time-negative RCFG and conventional CFG have almost identical inference times for a single denoising step case. However, as the number of denoising steps increases, the difference in inference time from conventional CFG to both Self-Negative and Onetime-Negative RCFG becomes more pronounced. At denoising step 5, a speed improvement of 2.05x is observed with Self-Negative RCFG and 1.79x with Onetime-Negative RCFG, compared to conventional CFG.
@@ 4.3 Ablation study
In our ablation study, as summarized in Table 3, we evaluate the average inference time of our proposed method under various configurations to understand the contribution of each component. Our proposed StreamDiffusion achieves an average inference time of 10.98/9.42 ms and 26.93/26.30 ms for denoising steps 1 and 4 on image-to-image/text-to-image generation, respectively. When the stream batch processing is removed (’w/o stream batch’), we observe a large time consumption increase, especially at 4 denoising steps. The absence of TensorRT (’w/o TRT’) leads to a more increase in time cost. The removal of pre-computation also results in increased time cost but not much. We attribute the reason to the limited number of key-value computations in StableDiffusion. Besides, the exclusion of input-output queue (’w/o IO queue’) also demonstrates an impact on average inference time, which mainly aims to optimize the parallelization issue results from pre- and post-processing. In the AutoPipelineImage2Image’s adding noise function, the precision of tensors is converted from fp32 to fp16 for each request, leading to a decrease in speed. In contrast, the StreamDiffusion pipeline standardizes the precision of variables and computational devices beforehand. It does not perform tensor precision conversion or computation device transfers during inference. Consequently, even without any optimization (’w/o any optimization’), our pipeline significantly outperforms the AutoPipelineImage2Image in terms of speed. We also evaluate the impact on the inference time of our pipeline with additional modules such as SSF and Self-Negative RCFG. As shown in Table. 2 and Table. 3, these additional modules, which only perform lightweight matrix and vector operations, do not have a noticeable effect on inference time.
4.4 Qualitative Results
@@
The generation results using the StreamDiffusion pipeline for real-time image-to-image transformation are presented in Fig. 1. This pipeline enables image generation with very low throughput from input images received in real time from cameras or screen capture devices. At the same time, it is capable of producing high quality images that effectively align to the specified prompt conditions. These capabilities demonstrate the applicability of our pipeline in various real time applications, such as real time game graphic rendering, generative camera effect filters, real time face conversion, and AI-assisted drawing.
@@
The alignment of generated images to prompt conditioning using Residual Classifier-Free Guidance (RCFG) is depicted in Fig. 10. The generated images, without using any form of CFG, exhibit weak alignment to the prompt, particularly in aspects like color changes or the addition of non-existent elements, which are not effectively implemented. In contrast, the use of CFG or RCFG enhances the ability to modify original images, such as changing hair color, adding body patterns, and even incorporating objects like glasses. Notably, the use of RCFG results in a stronger influence of the prompt compared to standard CFG. RCFG, although limited to image-to-image applications, can compute the vector for negative conditioning while continuously referencing the latent value of the input image and the initially sampled noise. This approach yields more consistent directions for the negative conditioning vector compared to the standard CFG, which uses UNet at every denoising step to calculate the negative conditioning vector. Consequently, this leads to more pronounced changes from the original image. However, there is a trade-off in terms of the stability of the generated results. While Self-Negative RCFG enhances the prompt’s effectiveness, it also has the drawback of increasing the contrast of the generated images. To address this, adjusting the $delta$ in Eq. 7 can modulate the magnitude of the virtual residual noise vector, thereby mitigating the rise in contrast. Additionally, using Onetime-Negative RCFG with appropriately chosen negative prompts can effectively mitigate contrast increases while improving prompt adherence, as observed in Fig. 10. This approach allows the generated images to blend more naturally with the original image.
Finally, the quality of standard text-to-image generation results is demonstrated in Fig. 11. Using the sd-turbo model, high-quality images like those shown in Fig. 11 can be generated in just one step. When images are produced using our proposed StreamDiffusion pipeline and sd-turbo model in an environment with GPU: RTX 4090, CPU: Core i9-13900K, and OS: Ubuntu 22.04.3 LTS, it’s feasible to generate such high-quality images at a rate exceeding 100fps. Furthermore, by increasing the batch size of images generated at once to 12, our pipeline can continuously produce approximately 150 images per second. The images enclosed in red frames shown Fig. 11 are generated in four steps using community models merged with LCM-LoRA. While these LCM models require more than 1 step for high quality image generation, resulting in a reduction of speed to around 40fps, these LCM-LoRA based models offer the flexibility of utilizing any base model, enabling the generation of images with diverse expressions.
@@ 5 Conclusion
We propose StreamDiffusion, a pipeline-level solution for interactive diffusion generation. StreamDiffusion consists of several optimization strategies for both throughput and GPU usage, including stream batch, residual classifier-free gudiance (RCFG), IO-queue for parallelization, stochastic similarity filter, pre-computation, Tiny AutoEncoder and the use of model acceleration tool. The synergistic combination of these elements results in a marked improvement in efficiency. Specifically, StreamDiffusion achieves up to 91.07 frames per second (fps) on a standard consumer-grade GPU for image generation tasks. This performance level is particularly beneficial for a variety of applications, including but not limited to the Metaverse, online video streaming, and broadcasting sectors. Furthermore, StreamDiffusion demonstrates a significant reduction in GPU power consumption, achieving at least a 1.99x decrease. This notable efficiency gain underscores StreamDiffusion’s potential for commercial application, offering a compelling solution for energy-conscious, high-performance computing environments.

### v2

@@ 3.4 Stochastic Similarity Filter
Figure 6: Average inference time comparison between Stream Batch and normal sequential denoising with using TensorRT.
Note: We emphasize that compared to determining whether we skip the compute via a hard threshold, the proposed probability-sampling-based similarity filtering strategy leads to a smoother video generation. Because the hard threshold is prone to making the video stuck, which hurts the impression of watching video streaming, while the sampling-based method significantly improves the smoothness. For the other efficiency improvement methods, we illustrate them in the supplementary material.
@@ 4 Experiments
We implement StreamDiffusion pipeline upon LCM, LCM-LoRA [25, 26] and SD-turbo [36]. As a model accelerator, we use TensorRT and for the lightweight efficient VAE, we use TAESD [17]. Our pipeline is compatible with the customer-level GPU. We test our pipeline on NVIDIA RTX4090 GPU, Intel Core i9-13900K CPU, Ubuntu 22.04.3 LTS, and NVIDIA RTX3060 GPU, Intel Core i7-12700K, Windows 11 for image generation. We note that we evaluate the throughput mainly via the average inference time per image through processing 100 images.
4.1 Quantitative Evaluation
We compare our method with the AutoPipelineForImage2Image, which is a pipeline developed by Huggingface diffusers  [^1] https://github.com/huggingface/diffusers. The average inference time comparison is presented in Table. 1. Our pipeline demonstrates a substantial speed increase. When we use TensorRT, StreamDiffusion achieves a minimum speed-up of 13.0 times when running the 10 denoising steps, and reaching up to 59.6 times in scenarios involving a single denoising step. Even though without TensorRT, StreamDiffusion achieves a 29.7 times speed up compared to AutoPipeline when using one step denoising, and an 8.3 times speedup at 10 step denoising.
Table 1: Comparison of average inference time (ms) at different denoising steps with speedup factors. The first column denotes the denoising steps and the AutoPipeline is from Diffusers [15].
Step StreamDiffusion StreamDiffusion w/o TRT AutoPipeline Img2Img
@@ 4.1 Quantitative Evaluation
Efficiency comparison regarding Stream Batch.
The efficiency comparison between Stream Batch and the original sequential U-Net loop is shown in Fig. 6. When implementing a denoising batch strategy, we observe a significant improvement in processing time. It achieves a reduction by half when compared to a conventional U-Net loop at sequential denoising steps. Even though applying TensorRT, the accelerator tool for neural modules, our proposed Stream Batch still boosts the efficiency of the original sequential diffusion pipeline by a large margin at different denoising steps.
Efficiency comparison regarding R-CFG.
Table. 2 presents a comparison of the inference times for StreamDiffusion pipelines with R-CFG and conventional CFG. The additional computations required to apply Self-Negative R-CFG are merely lightweight vector operations, resulting in negligible changes in inference time compared to when Self-Negative is not used. When employing Onetime-Negative R-CFG, additional UNet computations are necessary for the first step of the denoising process. Therefore, One-time-negative R-CFG and conventional CFG have almost identical inference times for a single denoising step case. However, as the number of denoising steps increases, the difference in inference time from conventional CFG to both Self-Negative and Onetime-Negative R-CFG becomes more pronounced. At denoising step 5, a speed improvement of 2.05x is observed with Self-Negative R-CFG and 1.79x with Onetime-Negative R-CFG, compared to conventional CFG.
Table 2: Comparison of average inference time (ms) at different denoising steps among different CFG methods
Step Self-Negative R-CFG Onetime-Negative R-CFG CFG
@@ 4.3 Ablation study
Figure 7: Ablation study on different components.
In our ablation study, as summarized in Fig. 7, we evaluate the average inference time of our proposed method under various configurations to understand the contribution of each component. Our proposed StreamDiffusion achieves an average inference time of 10.98/9.42 ms and 26.93/26.30 ms for denoising steps 1 and 4 on image-to-image/text-to-image generation, respectively. When the R-CFG is not used, we observe this results in the second largest efficiency drop. This demonstrates that R-CFG is one of the most critical components in our pipeline. When the Stream Batch processing is removed (’w/o stream batch’), we observe a large time consumption increase, especially at 4 denoising steps. We also evaluate the impact on the inference time of our pipeline regarding SSF. We observe that SSF plays a significant role in enabling energy saving and does not introduce extra time cost.
@@
Besides, the absence of TensorRT (’w/o TRT’) leads to a large increase in time cost. The removal of pre-computation also results in increased time cost but not much. We attribute the reason to the limited number of key-value computations in Stable Diffusion. Besides, the exclusion of input-output queue (’w/o IO queue’) also demonstrates an impact on average inference time, which mainly aims to optimize the parallelization issue resulting from pre- and post-processing. In the AutoPipelineImage2Image’s adding noise function, the precision of tensors is converted from fp32 to fp16 for each request, leading to a decrease in speed. In contrast, the StreamDiffusion pipeline standardizes the precision of variables and computational devices beforehand. It does not perform tensor precision conversion or computation device transfers during inference. Consequently, even without any optimization (’w/o any optimization’), our pipeline significantly outperforms the AutoPipelineImage2Image in terms of speed.
4.4 Quantitative Evaluation for the Image Quality.
We conduct a quantitative evaluation on the image quality. Specifically, we first evaluate the FID and CLIP scores on the text-to-image generation. We use the same dataset as [6] for the evaluation. We use LCM as our main baseline for the comparison. Note that our method is never trained; our method still improves LCM by a large margin in terms of the FID (29.69 vs. 26.79) and maintains a similar CLIP score (24.95 vs. 24.99). This demonstrates the effectiveness of our proposed method.
User study.
We also conduct a user study to validate the visual quality of different components of our StreamDiffusion. The results are shown in the Table. 3. Our Stream Batch with future-frame attention significantly enhances time consistency compared to its absence. Additionally, our SSF method addresses the crucial yet rarely explored issue of energy efficiency. While SSF is simple, it is far from trivial: a common approach might apply a hard threshold on similarity to regulate streaming flow. However, as noted in the main text, we introduce a novel approach—using probability sampling to achieve superior streaming quality. As Table 3 illustrates, this approach is preferred by more users over the hard-threshold method for visual quality. Moreover, compared to the vanilla CFG method, both our self-Negative R-CFG and one-time R-CFG are more preferred by the users, demonstrating our method not only improves efficiency but also visual quality.
# Users Without Future-Frame (%) With Future-Frame (%)
486 9.67 90.33
# Users Without SSF (%) With SSF (%)
144 48.61 51.39
# Users Hard Threshold (%) SSF (%)
45 24.44 75.56
# Users With / Without CFG (%) Self-Neg R-CFG / One-Time R-CFG (%)
257 22 / 20 35 / 23
Table 3: User study results: Future-frame attention consistency (486 users), SSF quality imperceptibility (144 users), streaming quality between hard threshold filter and SSF (45 users), comparison of with/without CFG and self-negative/one-time R-CFG (257 users).
@@
This pipeline enables image generation with very low throughput from input images received in real-time from cameras or screen capture devices. At the same time, it is capable of producing high-quality images that effectively align to the specified prompt conditions. These capabilities demonstrate the applicability of our pipeline in various real-time applications, such as real-time game graphic rendering, generative camera effect filters, real-time face conversion, and AI-assisted drawing.
@@
The alignment of generated images to prompt conditioning using Residual Classifier-Free Guidance (R-CFG) is depicted in Fig. 8. The generated images, without using any form of CFG, exhibit weak alignment to the prompt, particularly in aspects like color changes or the addition of non-existent elements, which are not effectively implemented. In contrast, the use of CFG or R-CFG enhances the ability to modify original images, such as changing hair color, adding body patterns, and even incorporating objects like glasses. Notably, the use of R-CFG results in a stronger influence of the prompt compared to standard CFG. R-CFG, although limited to image-to-image applications, can compute the vector for negative conditioning while continuously referencing the latent value of the input image and the initially sampled noise. This approach yields more consistent directions for the negative conditioning vector compared to the standard CFG, which uses UNet at every denoising step to calculate the negative conditioning vector. Consequently, this leads to more pronounced changes from the original image. However, there is a trade-off in terms of the stability of the generated results. While Self-Negative R-CFG enhances the prompt’s effectiveness, it also has the drawback of increasing the contrast of the generated images. To address this, adjusting the $delta$ in Eq. 6 can modulate the magnitude of the virtual residual noise vector, thereby mitigating the rise in contrast. Additionally, using Onetime-Negative R-CFG with appropriately chosen negative prompts can mitigate contrast increases while improving prompt adherence, as observed in Fig. 8. This approach allows the generated images to blend more naturally with the original image.
Besides, Fig. 9 in appendix shows the image-to-image generation results using StreamBatch Cross-frame attention, with 4 denoising steps. As evident from the figure, compared to the results of StreamDiffusion without Cross-frame attention, the method incorporating information from future and past frames exhibits increased temporal consistency.
@@ 5 Conclusion
We propose StreamDiffusion, a pipeline-level solution for interactive diffusion generation. StreamDiffusion consists of several optimization strategies for both throughput and GPU usage, including Stream Batch with cross-attention, residual classifier-free guidance (R-CFG), IO-queue for parallelization, stochastic similarity filter, pre-computation, Tiny AutoEncoder, and the use of the model acceleration tool. The synergistic combination of these elements results in a marked improvement in efficiency. Specifically, StreamDiffusion achieves up to 91.07 frames per second (fps) on a standard consumer-grade GPU for image generation tasks. This performance level is particularly beneficial for a variety of applications, including but not limited to the Metaverse, online video streaming, and broadcasting sectors. Furthermore, StreamDiffusion demonstrates a significant reduction in GPU power consumption, achieving at least a 1.99x decrease. This notable efficiency gain underscores StreamDiffusion’s potential for commercial application, offering a compelling solution for energy-conscious, high-performance computing environments.
@@ 6 Acknowledgments
Appendix A More Architecture details
A.1 Input-Output Queue
A.2 Pre-computation
The U-Net architecture requires both input latent variables and conditioning embeddings. Typically, the conditioning embedding is derived from a text prompt, which remains constant across different frames. To optimize this, we pre-compute the prompt embedding and store it in a cache. In interactive or streaming mode, this pre-computed prompt embedding cache is recalled. Within U-Net, the Key and Value are computed based on this pre-computed prompt embedding for each frame. We have modified the U-Net to store these Key and Value pairs, allowing them to be reused. Whenever the input prompt is updated, we recompute and update these Key and Value pairs inside U-Net.
For consistent input frames across different timesteps and to improve computational efficiency, we pre-sample Gaussian noise for each denoising step and store it in the cache. This approach is particularly relevant for image-to-image tasks.
$x_{t}=\sqrt{\alpha_{\tau}}x_{0}+\sqrt{\beta_{\tau}}\epsilon$ (10)
This is a minor point in low throughput scenarios, but at frame rates higher than 60 FPS, the overhead of recomputing these static values becomes noticeable.
We note that we have a specific design for the inference parameterization for latent consistency models (LCM). As per the original paper, we need to compute $c_{\mathrm{skip}}(\tau)$ and $c_{\mathrm{out}}(\tau)$ to satisfy the following equation:
A.3 Model Acceleration and Tiny AutoEncoder
Appendix B Text-to-Image Quality
The quality of standard text-to-image generation results is demonstrated in Fig. 11. Using the sd-turbo model, high-quality images like those shown in Fig. 11 can be generated in just one step. When images are produced using our proposed StreamDiffusion pipeline and SD-turbo model in an environment with GPU: RTX 4090, CPU: Core i9-13900K, and OS: Ubuntu 22.04.3 LTS, it’s feasible to generate such high-quality images at a rate exceeding 100fps. Furthermore, by increasing the batch size of images generated at once to 12, our pipeline can continuously produce approximately 150 images per second. The images enclosed in red frames shown in Fig. 11 are generated in four steps using community models merged with LCM-LoRA. While these LCM models require more than 1 step for high quality image generation, resulting in a reduction of speed to around 40fps, these LCM-LoRA based models offer the flexibility of utilizing any base model, enabling the generation of images with diverse expressions.
Appendix C GPU Usage Under Dynamic Scene
We also evaluate the GPU usage under dynamic scenes on one RTX 4090 GPU, as shown in the Figure. 13. The analysis of the GPU usage is shown in Section 4.2 of the main text.
Figure 12: GPU Usage comparison under static scene. (GPU: RTX3060, Number of frames: 20) The blue line represents the GPU usage with SSF, the orange line indicates GPU usage without SSF, and the red line denotes the Skip probability calculated based on the cosine similarity between input frames. Additionally, the top of the plot displays input images corresponding to the same timestamps. In this case, the character in the input images is only blinking.